OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 3x + 15y = 33 4x − 3y = 6

OpenStudy (anonymous):

x = 33 - 15y / 3

sam (.sam.):

Isolating x, 3x + 15y = 33 3x=33-15y \[x=\frac{33-15y}{3}\] \[x=11-5y\] substitute to second equation 4x − 3y = 6 4(11-5y) − 3y = 6 solve for y